OpenStudy (anonymous):

Will give medal help. Choose one Addition property of equality , Subtraction property of equality, Multiplication property of equality, Division property of equality, Simplifying, Distributive property

OpenStudy (anonymous):

where is the problem?

OpenStudy (anonymous):

\[\frac{ 7(v+2) }{7 }=\frac{ -35 }{ 7 }\]

OpenStudy (anonymous):

\[v+2=-5\]

OpenStudy (anonymous):

v+2-2=-5-2

OpenStudy (anonymous):

v=-7

OpenStudy (anonymous):

1.) since 7(v+2) means distribute 7v + 14, so distrubitive property 2.) division property 3.) subtraction property

OpenStudy (anonymous):

you are right i plug it in and got 5=5

OpenStudy (anonymous):

the 5 are -

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

@n648c788 what about the v=-7

OpenStudy (anonymous):

the would be additive property, but you can also divide

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

np, feel free to notify me for help if needed